OpenStudy (anonymous):

To solve for the following proportion, you could use equivalent ratios, divide out the known ratio, or use cross products. Choose one of these methods to solve for t. Show your work and explain why you chose that method. 18over5 = 4overT

OpenStudy (anonymous):

i would suggest cross multiplying

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

18/5=4/t 18t=20 t=1.11

OpenStudy (anonymous):

(18/5)=(4/t)=18t=20=t=20/18

OpenStudy (anonymous):

or 10/9 in simplest form